The Harvard Family of Bibliography Styles

Peter Williams (peterw@archsci.arch.su.edu.au)

April 15, 1994

1 Introduction

This document describes the harvard family of bibliographic styles which are provided in addition to those described in Lamport (1986) and Patashnik (1988). This style is primarily intended for use with the BibTEX bibliographic database management system. However, provision is also made for hand coding of bibliographies.

2 Citations

There are two primary forms of citation in the harvard style dependent upon whether the reference is used as a noun or parenthetically. Additionally, where there are more than two authors, all authors are listed in the first citation and in subsequent citations just the first author’s name followed by ‘et al.’ is used. The following example from Pitson (1978) illustrates these points.

The major improvement concerns the structure of the interview (Ulrich & Trumbo 1965, p. 112) . . .. Later reports (Carlson, Thayer, Mayfield & Peterson 1971) record greatly increased interviewer reliability for structured interviews. Wright (1969, p. 408) comments that ‘undoubtedly interviewer skill is directly related to the validity, quantity and quality of the interview output’, and this would suggest some sort of interviewer training is called for. Rowe (1960), for example, found that trained interviewers are better able to evaluate applicants with some measure of reliability. In addition Wexley, Sanders & Yukl (1973) showed that by extensive interviewer training all significant contrast effects could be eliminated. The results of the 1971 study (Carlson et al. 1971) are still relevant, but efforts to . . ..

To facilitate using a citation as a noun a new command \citeasnoun has been created which has the same syntax as the \cite command except that multiple citations are not permitted. The effect of this command is that

As \citeasnoun{btxdoc} and \citeasnoun[Annex~B]{latex} describe \ldots

produces

As Patashnik (1988) and Lamport (1986, Annex B) describe …

whereas

The \BibTeX\ \cite{btxdoc} and \LaTeX\ \cite[Annex~B]{latex} manuals \ldots

produces

The BibTEX (Patashnik 1988) and LaTEX (Lamport 1986, Annex B)

manuals …

A second new command \possessivecite is provided for those instances where it is desired to use the citation as a possessive noun phrase. This is a variation on the \citeasnoun command and multiple citations are not permitted. As an example of its use

\possessivecite{latex} description of this feature is \ldots

produces

Lamport’s (1986) description of this feature is …

A third new command \citeaffixed allows text to be affixed inside the beginning of the parenthesis of a parenthetical citation. This command is like the the \cite command except that it takes a second argument – the text to be affixed after the opening parenthesis. For example

\BibTeX\ manuals \citeaffixed{latex,btxdoc}{e.g.} describe \ldots

produces

BibTEX manuals (e.g. Lamport 1986, Patashnik 1988) describe …

By default, where appropriate, citations are abbreviated automatically after the first reference when bibliographies are produced by BibTEX. Provision is also made for this feature to be accessed during manual coding. This feature may be overridden by using the \citationmode command which takes full, abbr or default as its single argument. The command \citationmode{full} makes the system use full citations, \citationmode{abbr} makes the system use abbreviated citations and \citationmode{default} causes the default behaviour of using full citations for the first instance and abbreviated citations thereafter.

In addition to these primary forms of citation, the citation commands \citeyear and \citename are provided. \citeyear behaves like the \cite command except that only the year portion of the citation label is used. For example,

\citeyear{btxdoc,latex}

produces (1988, 1986). \citename behaves like the \citeasnoun command except that only the author name(s) (unabbreviated) portion of the citation label is used. For example,

\citename{btxdoc}

produces

Patashnik.

The use of this command does not trigger the use of abbreviated citations for subsequent \citeasnoun and \cite references.

3 Styles

3.1 Bibliography Styles

There are four bibliography styles currently available within the harvard family, agsm (used in this document) which is based on Pitson (1978, pp. 95–98), dcu which is based upon the conventions in use in the Design Computing Unit, Department of Architectural and Design Science, University of Sydney, kluwer which aspires to conform to the requirements of Kluwer Academic Publishers and nederlands which conforms to Dutch conventions. They are invoked by the \bibliographystyle as described in Lamport (1986, p. 74) and effect the layout of the entries in the bibliography.

3.2 Citation Styles

There are two citation styles currently available within the harvard family, agsm (used in this document) and dcu which for the previous example would produce:

The BibTEX (Patashnik, 1988) and LaTEX (Lamport, 1986, Annex B) manuals …

and for multiple citations such as

  The original documentation \cite{btxdoc,latex} say \ldots

the agsm citation style produces

The original documentation (Patashnik 1988, Lamport 1986) say …

and the dcu citation style produces

The original documentation (Patashnik, 1988; Lamport, 1986) say … The default citation style is agsm and both styles have no effect on the appearance of the \citeasnoun citation format.

These styles are invoked by the \citationstyle command, for example:

  \citationstyle{agsm}.

Because these styles affect the format of parenthetical citations, this command should appear before any \cite commands.

3.3 Parenthesis Style

The type of parenthesis used in citations may be set using the \harvardparenthesis command which takes one argument. The argument to this command must be one of round, curly, angle or square. The default value is round. If the bibliographic style chosen is agsm or dcu then the parenthesis style chosen using \harvardyearparenthesis is used with the year portion of the entries in the bibliographic listing.

Authors of style files for use with the harvard family who wish to make use of this feature should use the strings " \harvardleft " and " \harvardright " instead of the respective parenthesis characters where they wish them to be effected by the selection made with \harvardparenthesis.

3.4 Conjunction Style

In the previous examples for the agsm bibliographic style a “&” character is used to signify conjunction between a pair of names or between the last two names of a list of names. Similarly the word “and” is used for the dcu style. With these two styles this convention may be overwritten by using \renewcommand to redefine the command \harvardand. This should be done after the \citationstyle command (if used) as this command resets it to the default for the style selected.

4 Doing It By Hand

Hand coding is accomplished much the same as described in Lamport (1986, p. 73) except that the new command \harvarditem is used in place of \bibitem. The syntax of this command is

\harvarditem [abbr-citation]{full-citation}{citation-year}{cite-key}

where

abbr-citation is the (optional) abbreviated citation (minus the year) to be used in the text subsequent to the first mention of a particular reference, full-citation is the full citation (minus the year) to be used in the text on the

first mention of a particular reference,

citation-year the year portion of the citation including any suffices required to disambiguate citations, and

cite-key is the key used in the \cite and \citeasnoun commands.

5 Acknowledgement

The motivation for this style came from Fay Sudweeks of the Design Computing Unit who also originated the formats for the dcu style and proofread their implementation.

The nederlands bibliographic style was implemented by Werenfried Spit (spit@vm.ci.uv.es).

The idea for \citeyear came from Renate Schmidt (Renate.Schmidt@mpisb.mpg.de).

The solution to the mysterious \enddocument problem came from Berwin A. Turlach (berwin@core.ucl.ac.be).

References

Lamport, L. (1986), LaTEX: A Document Preparation System, Addison-Wesley.

Patashnik, O. (1988), BibTEXing, Documentation for general BibTEX users.

Pitson, J. (1978), Style Manual for authors editors and printers of Australian government publications, 3rd edn, Australian Government Publishing Service, Canberra.